# Human-Machine Interfaces for Subsea Telerobotics: From Soda-straw to Natural Language Interactions

### v1

@@ Abstract
NL-based interfaces leverage the natural human way of communicating and reasoning for more intuitive mission planning and control. The process began as experimental research almost two decades ago [112, 113], aiming to simplify human-robot communication by enabling operators to issue high-level commands using everyday speech rather than complex control inputs. Classic visual language systems such as RoboChatTag [114] and RoboChatGest [115] use sequences of symbolic patterns to convey simple instructions from diver to robot/operator, utilizing AR-Tag markers and hand gestures, respectively. Recently, natural language processing (NLP) with deep learning networks has allowed the smooth conversion of spoken/written language into understandable and actionable commands for robotic systems [116]. Early NL interfaces focus on basic navigation commands or status queries such as “take a left", “report battery status" [117], and have since evolved to support more sophisticated mission planning and contextual understanding. Fig. 4 provides a chronological overview of NL integration in telerobotic control and mission management.
@@
Figure 4: A timeline progression of natural language integration in subsea telerobotic interfaces and mission management tasks over the past $1515$ years is shown. The key works referenced are: Hallin et al. [113], PUMMA [118], REGIME [119], Robb et al. [40], RoboChat [116], OceanChat [120], ChatSim [121], OceanGPT [122], OceanPlan [48], and Word2Wave [43].
More recently, the success of large language models (LLMs) has led roboticists to utilize detailed language descriptions to plan robotic missions. LLM-driven mission planning systems reported in contemporary literature are OceanChat [120], Word2Wave [43], and OceanPlan [48]. These interfaces operate as offline planners, processing full mission descriptions and planning the optimum trajectory before deployment [118]. The planner first converts spoken or written instructions into structured commands using speech recognition and syntactic parsing. This is followed by semantic analysis, where the system interprets the meaning and maps it to specific robotic actions or a sequence of tasks. A particular example for Word2Wave interface [43] is shown in Fig. 5. When an operator specifies: “spiral for $55$ turns to a radius of $5050$ m at a speed of $11$ m/s”; the system breaks down the request into individual sub-tasks and plans accordingly such as setting velocity, setting altitude, and finally making spiral move to reach the specified radius. Such interfaces enable a robotic system to plan and execute complex missions from high-level operator instructions.
In contrast to offline planning interfaces, online interfaces allow real-time dialogue between the operator and
@@
Figure 5: An example of a language-guided mission planning interface [43]. The robot parses verbal instructions into sub-tasks such as S (start), Mv (move), Sp (spiral), Az (azimuth), E (end); and then generates mission waypoints.
@@
the robot. Hastie et al. [119] demonstrate an online dialogue module named REGIME (Report Generation from Metadata), where sensor data and event logs are summarized into mission updates upon operator demand. Robb et al. [40] introduce a chat-based interface for controlling a group of robots, allowing the operator to inquire about the status of individual robots via the leader and adjust their behavior as needed. This reduces the cognitive load on operators by shifting focus from managing multiple robots to a single control point through the leader. In ChatSim [121], dialogue capabilities are utilized to customize a simulated environment. The operator first asks for specific structures (e.g. canyon, shipwreck, coral reef, oyster) to be added into the virtual world, followed by issuing commands to the robot such as “go through the canyon and find the shipwreck”, “move in a circle of radius $33$ m over the oysters”, etc.
The current challenges with integrating language models into telerobotic systems are multifaceted. First, the computational demands of LLMs may exceed the capabilities of resource-constrained underwater systems. To address this, researchers are developing smaller language models (SLMs) while technologies such as LoRA [123] are helping to adapt LLMs for real-time applications in edge devices. Second, ensuring that the robot accurately interprets commands and understands spatial references requires sophisticated models, extensive training data, and robust error-handling mechanisms [116]. The consequences of misinterpretation can be severe in high-risk subsea operations, hence current research focuses on making these systems more resilient, context-aware, and capable of handling complex, multi-step instructions [124]. Moreover, online commands can be error-prone due to variations in dialect, tone, or pronunciation [125]. Instead, an offline planning interface allows the operator to review and correct for potential errors before launch. Further discussions on this are presented in Sec. 6.
@@
Figure 6: A digital twin of an underwater ROV is shown in action; (a) a human is teleoperating it to pick up an object using head movements and finger grips; (b) the twin model rendered in HoloOcean is mimicking the action in virtual environment.
8 3 Simulators and Digital Twins
@@
Simulation environments offer a valuable alternative for prototyping, error debugging, and pilot training for marine robotic systems in a controlled, risk-free setting, without requiring expensive logistics and access to real waterbodies. These simulators not only replicate the dynamics of underwater conditions but also support the integration of digital replicas of physical systems, facilitating the testing and development of robust control algorithms and mission planning strategies for real robots. Fig. 6 shows an example of underwater grasping with a digital twin rendered in the HoloOcean simulator.
9 3.1  Digital Twins (DTs)
@@
A digital twin refers to a simulated replica of a real robot that possesses similar attributes and can mimic the motion of its real-world pair within a virtual environment [126, 127]. The concept of DT for field robots was first presented by NASA, where the digital entity was defined as an integrated multi-physics, multi-scale, probabilistic simulation of a vehicle or system that mirrors the life of its real-world peer [128]. This project envisioned four major applications of the DT technology.
  • Pre-mission simulation: The twin will simulate the vehicle’s future missions before launch, enabling studies of mission parameters, anomalies, and mitigation strategies to enhance mission success.
  • Real-time flight mirroring: The twin will continuously update and mirror the actual flight with real-time sensor data, predicting flight status for ongoing analysis and adjustment.
  • In-situ forensics: The twin will allow diagnosis and analysis of potential catastrophic faults in real-time and address anomalies during the flight.
  • Mission modification analysis: The twin will assess the impact of unforeseen mission parameter changes (e.g., failed actuators), and analyze new load distributions to inform mission decisions.
DTs exhibit varying levels of complexity, ranging from basic motion replication to comprehensive emulation of physical and operational attributes. The simplified setup shown in Fig. 6 does not replicate the real testbed in the simulator; however, the digital robot mimics the motion of the real ROV, thereby qualifying as its twin. For more sophisticated DTs, the digital replica must mirror key physical properties (e.g. dimensions, weight, buoyancy), energy consumption profile, sensor configurations, and interaction with the external environment. Examples of some SOTA simulators and their DT collections include BlueROV2 in UNav-Sim [129], Triton XLX in PaleBlue [130] and VMAX simulator [131], VideoRay Pro5 and Defender in Era’s VideoRay simulator [132], and Girona500 in UWSim [133]; see Fig. 7.
@@
The use of DTs has been reported in different branches of robotics including aerial, manufacturing, surgical, space, and marine robots. All DT systems are built upon some common backends with domain-specific techniques for data acquisition and rendering [127]. Generally, the raw sensor data acquired from the robot is processed with physics engines and reconstructed as a high-fidelity DT model [127]. Considering the complex nature of subsea environments, Van et al. [32] identify the communication between the two entities as a fundamental component of a successful subsea DT system. The authors highlight three use cases for DT: (i) ROV launch and recovery training, (ii) failure prediction and remaining useful life estimation, and (iii) trust-aware DT development to reduce operator’s decision-making time in critical situations. However, other literature [30, 135] argue that DT systems do not require a fully reliable or accelerated communication mode. Instead, during a communication disruption, the predictive nature of a DT system can be utilized to create a realistic representation as if communication had remained uninterrupted. DT systems have also been deployed in underwater sensor networks to test communication reliability and synchronization accuracy between multiple nodes [136]. More recently, Adetunji et al. [29] integrate SLAM and path planning capabilities in a versatile DT system. Other applications of DT in marine robotics include pipeline inspection [137], assistive control [13], trajectory planning [30], etc.
10 3.2  Underwater Simulators
@@
A simulator is essential to work with DTs as well as to mimic the external environment parameters. However, given the complex fluid dynamics and other physicochemical variations, it is extremely difficult to design an accurate physics-based model of an underwater environment. Engineers have created aquatic environments with physics and game engines (e.g. Unity, Blender, WebGL3D, ODE) as backends and have integrated basic properties of the medium such as buoyancy, waves, current profile, etc.; examples include UUV Simulator [26] and UWSim [133] that support Gazebo plugin [138] as well. Independently designed virtual ROVs or DTs of real ROVs are rendered in these platforms to test their kinematic properties and interactions with the surroundings. Researchers have studied different underwater vehicles such as gliders [30, 139, 140], ROVs [135, 26], AUVs [137], underwater legged robots [141], and humanoid robots [111] in these simulators. Yang et al. [30] investigate a glider trajectory in a deep virtual ocean. Their simulations show the effects of the current gradient across the water column on the glider’s trajectory, especially during the vertical dives. Hu et al. [139] model a blended-wing-body underwater glider (BWBUG) and investigate the accuracy of a pitch controller under the influence of ocean currents.
@@
Table 3: Comparison of underwater simulators are shown. Basic state estimation sensor plugins (e.g., IMU, GPS, depth) are common for all simulators, therefore omitted here; indicates digital twin of a real ROV.
@@
  Sensor Plugin Built-in Scenarios Robot Models
  PaleBlue [130] RGB-D camera, laser Reefs, oil and gas rigs Triton XLX
Abyssal [134] Camera Rigs, sand cloud, pipelines Abyssal ROV
M8-R Sim [142] Camera 140+ coral reefs with starfishes M8-R ROV
Era VideoRay [132] Camera, sonar, lights 6 seabed scenarios with shipwrecks VideoRay Pro5 , Defender
GRi VROV [67] Camera and sonar Pipelines and rigs LittleGeek, BigGeek
UUV Simulator [26] Camera, sonar, DVL, CPC Lakes, shipwrecks, coast, BOP panels A9 , LAUV , Saga , RexROV
UUVSim [143] Camera, sonar Open water, current, seabed 6T UUV, RexROV
HoloOcean [27] Camera, sonar, beacon, DVL Dams, pier harbors, pipelines, open water 6 ROV agents
UWSim [133] Camera, DVL, multi-beam sonar Dredging, pipelines, seabed with amphora Girona500
UNav-Sim [129] Camera, LiDAR Seabed with algae and sand-covered pipelines BlueROV2 heavy
MARUS [144] Camera, LiDAR, sonar, DVL, AIS Ships, human divers, seabed One AUV
StoneFish [145] RGB-D camera, sonar, DVL Sea surface, seabed, wind, sun Girona500
NaviSuite [56] Camera, sonar, DVL, USBL Cables, pipelines, and ship anchors VideoRay Defender
VMAX Simulator [131] Camera, sonar Dynamic seabed, cables, ship with crane Triton , Comanche , Mojave
ACSM [146] Camera Rigs and pipelines Triton XLX
Unicom [147] FPV and TPV cameras Pipelines and rigs with valves, knobs, doors 2 work class ROVs
@@
Fig. 7 illustrates some SOTA underwater simulators developed across academia and industry. Among proprietary simulators, Abyssal [134] and ACSM [146] offer a rich collection of underwater infrastructures and work-class ROV models, designed to train operators in tasks such as underwater mining, pipeline repairing, and maintenance of offshore oil and gas rigs. PaleBlue [130] and VMAX Simulator [131] allow users to import custom-designed ROVs from CAD models and simulate ROV behavior, forces, frictions, etc. in the virtual world. Era Marine and EIVA have partnered with VideoRay to create advanced simulators for the VideoRay Pro5 and Defender ROVs. These simulators, Era’s VideoRay [132] and EIVA NaviSuite [56], provide advanced tools for planning, surveying, and underwater inspections. M8-R Sim [142], originally developed for the MATE ROV Competition, has over $140140$ simulation scenarios, including those for monitoring and estimating the growth of Crown of Thorns Starfish (COTS). Unlike higher-end professional simulators, open-source alternatives offer fewer built-in scenarios and ROV models but allow for better customization and further development according to mission-specific needs. For instance, UWSim [133], StoneFish [145], and HoloOcean [27] support numerous sensors e.g., side scan sonar, single beam and multi-beam imaging sonar, acoustic beacon, DVL, magnetometer, force sensor, etc. UUV Simulator [26] integrates chemical particles and a simulation scenario to measure their concentration using a CPC (condensation particle counter) sensor. They also provide closed water (lake) and open water (coast) scenarios, each with different current profiles. As shown in Table 3, the open-source solutions are better suited for research and development with ROS support, whereas licensed products offer superior training opportunities for commercial ROV operation.
11 4 Human-Machine Shared Autonomy
@@
Telerobotic systems are categorized into three types according to their intelligent decision-making capabilities: (i) direct teleoperator control with no intelligence, (ii) shared/collaborative reasoning, and (iii) autonomous/self-sufficient systems. In direct teleoperation, the operator has total control over the robot’s motion and actions, generally through a live feed and a motion controller. However, these methods are constrained by communication latency and operator workload, making them less efficient for long-term complex tasks. Shared autonomy features bridge the gap between the human and the machine by allowing robots to perform certain tasks autonomously while the operator manages high-level decisions [35, 22]. This reduces the operator’s cognitive load, compensates for communication delays, and enhances operation safety. In contrast, AUVs function independently with minimal human intervention [148]. These robots rely on advanced algorithms for perception, decision-making, and execution, making them ideal for repetitive tasks in familiar environments. However, their utility is constrained by environmental uncertainties, the cost of logistics, and the limitations of current AI-based technologies [149, 150].
12 4.1  Direct Teleoperation Mode
@@
Direct or hands-on teleoperation involves controlling the movements of an underwater robot using low-level commands without intelligent decision-making by the robot. The advantage of such systems is that the operator can execute instant decisions during the operation. It allows for responsive control, enabling precise adjustments that automated systems may struggle with. Despite high cognitive demand, direct teleoperation remains the preferred approach for manipulation tasks involving high risk and fine precision, where the operator controls one joint angle at a time in a joint-by-joint fashion using a traditional console [151]; examples include wreck inspection [152], valve control [153], benthic sampling [151], etc.
@@
These systems are challenged by the inherent communication delays in underwater environments, which can reduce control precision and responsiveness. The delays further complicate tasks that depend on real-time feedback [154, 153]. To address this, researchers have developed adaptive control strategies to compensate for both model uncertainties and external disturbances during teleoperation [155, 153]. Furthermore, electromagnetic and optical communication modes offer promising solutions for establishing high-speed reliable links [156, 157], improving the real-time commands and feedback flow in direct teleoperation systems.
@@
Table 4: The table summarizes various types of intelligent machines and their autonomous task-sharing capabilities reported across the subsea telerobotics literature.
@@
collaboration leader ROV follower ROVs search and rescue [6, 162]
@@
demonstration relevant tasks adapt to novel tasks sample collection [104, 165]
@@
Context-aware Provide contextual Medium Search for relevant features, Inspection, surveillance, [5, 166]
exploration cues or guidance report key findings environmental monitoring [167]
Docking Guide the robot Low Identify docking point, Charging, data transfer, [168, 169]
assistance close to the dock self-correct fine motion repair, recovery [170, 171]
13 4.2   Human-Robot Collaborative Reasoning
@@
The goal of collaborative decision-making systems is to offer a blend of human oversight along with automated functions. These systems allow operators to set high-level tasks or objectives while the robot handles low-level operations, thus reducing cognitive load on the operator and enabling more efficient operation [22]. For instance, an operator can command the simple autonomous task of hovering; the robot determines its state using an IMU, DVL, and depth sensors to correct for any drift motion to remain stationary. More advanced reasonings are achieved in a supervisory manner (for a single robot) or with a leader-follower strategy (for a group of robots). Table 4 provides an overview of the SOTA intelligent capabilities of underwater robots and their task-sharing mechanisms with humans.
@@
Mission planning, online mission adjustment, and fault diagnosis systems benefit from shared autonomy while keeping the operator in the loop [172], but without demanding continuous oversight. Research in supervised planning and navigation demonstrates that a new mission path suggested by a circumstance-aware robot is often safer and reduces task completion time [34, 20, 48, 173]. The GALILEO planning interface adjusts the course of actions and resource allocations based on high-level goals set by the operator [174]. Additionally, intelligent robots are capable of detecting and reporting internal faults [175] and data anomalies [159], reducing the need for constant operator oversight. Such distributed control also allows the operator to focus on mission objectives while offloading low-level tasks to the robot. Sharing scientific objectives with a context-aware robot is beneficial too, as the robot can search for specific cues or features under operator guidance. The CView project [5] provides an example; the operator defines the object of interest (e.g., wall, vessel, sluice) and the features to detect (e.g., cracks), enabling the robot to survey the objects and report key findings.
@@
SUPERMAN, the first supervisory subsea teleoperation interface, was proposed by NASA in $19781978$ [176]; however, the development has been slow due to the difficulty of designing high-level autonomy considering the environmental challenges. SAUVIM (1998199819981998[177] is one of the earliest semi-autonomous vehicles with a 6-DOF manipulator arm developed for deep water intervention tasks. More recently, supervisory control frameworks such as DexROV [109] and SHARC [21] have been developed where the operator performs remote manipulation tasks using language commands and hand gestures while experiencing the task in immersive virtual reality [34]. The GRASPER project [178] utilizes a pre-scanned scene point cloud to inform the robot about obstacle locations. The autonomous action is rendered in the UWSim simulator [133] in real-time, allowing the user to supervise grasping tasks in a virtual world.
@@
Moreover, contemporary researchers have developed adaptive supervision systems where the robot understands the operator’s overarching intention and adjusts the low-level reactions accordingly. Lee et al. [20] propose a shared teleoperation method where humans can provide real-time input by sketching a desired path on the visual console, guiding the intelligent system, which plans an optimum path along that drawn trajectory and executes the motion. The adaptive supervision system built within the GreenSea interface provides operators customized assistance based on their skill levels [35]. This helps novice pilots achieve performance comparable to that of experts.

### v2

@@ Abstract
NL-based interfaces leverage the natural human way of communicating and reasoning for more intuitive mission planning and control. The process began as experimental research almost two decades ago [82, 83], aiming to simplify human-robot communication by enabling operators to issue high-level commands using everyday speech rather than complex control inputs. Classic visual language systems such as RoboChatTag [84] and RoboChatGest [85] use sequences of symbolic patterns to convey simple instructions from diver to robot/operator, utilizing AR-Tag markers and hand gestures, respectively. Recently, natural language processing (NLP) with deep learning networks has allowed the smooth conversion of spoken/written language into actionable commands for robotic systems [86]. Early NL interfaces focus on basic navigation commands or status queries such as “take a left", “report battery status" [87], and have since evolved to support more advanced contextual understanding. Fig. 4 provides a chronological overview of NL functionalities in telerobotic control and mission management.
@@
Figure 4: Progression of natural language functionalities in subsea telerobotic interfaces and mission management tasks over the past $1515$ years is shown. The key works referenced are: Hallin et al. [83], PUMMA [88], REGIME [89], Robb et al. [90], RoboChat [86], OceanChat [91], ChatSim [92], OceanGPT [93], OceanPlan [31], and Word2Wave [94].
More recently, the success of large language models (LLMs) has led roboticists to utilize detailed language descriptions to plan robotic missions. LLM-driven mission planning systems reported in contemporary literature are OceanChat [91], Word2Wave [94], and OceanPlan [31]. These interfaces operate as offline planners, processing full mission descriptions and planning the optimum trajectory before deployment [88].
@@
Figure 5: An example of a language-guided mission planning interface is shown [94]; verbal instructions are parsed into sub-tasks such as S (start), Mv (move), Sp (spiral), Az (adjust), E (end) – to generate deployable mission waypoints.
The planner first converts spoken or written instructions into structured commands using speech recognition and syntactic parsing. This is followed by semantic analysis, where the system interprets the meaning and maps it to specific robotic actions. A particular example for Word2Wave interface [94] is shown in Fig. 5. When an operator specifies: “spiral for $55$ turns to a radius of $5050$ m at a speed of $11$ m/s”; the system breaks down the request into individual sub-tasks and plans accordingly such as setting velocity, setting altitude, and finally making spiral move to reach the specified radius. Such interfaces enable a robotic system to plan and execute complex missions from high-level operator instructions.
@@
In contrast to offline planning interfaces, online interfaces allow real-time dialogue between the operator and the robot. Hastie et al. [89] demonstrate REGIME dialogue where sensor data and event logs are summarized into mission updates upon operator demand. Robb et al. [90] introduce a chat-based interface for controlling a group of robots, allowing the operator to inquire about the status of individual robots via the leader and adjust their behavior as needed. This reduces the operator’s cognitive load by shifting focus from managing multiple robots to a single control point through the leader. In ChatSim [92], dialogue capabilities are utilized to customize a simulated environment. The operator first asks for specific structures (e.g. canyon, shipwreck, coral reef, oyster) to be added into the virtual world, followed by issuing navigation commands such as “go through the canyon and find the shipwreck”. However, online commands can be error-prone due to variations in dialect, tone, or pronunciation [95]. Instead, an offline planning interface allows the operator to review and correct potential errors before launch. Further discussions on this are presented in Sec. 7.1.
@@
Figure 6: A digital shadow of an underwater ROV is shown in action; (a) a human is teleoperating it to pick up an object using head movements and finger grips; (b) the replica rendered in HoloOcean is mimicking the action in virtual environment.
4 4. Simulators and Digital Twins
@@
Simulators offer a valuable alternative for prototyping, error debugging, and pilot training for marine robotic systems in a controlled, risk-free setting, without requiring expensive logistics and access to real waterbodies. They not only replicate the environmental dynamics but also support the integration of digital replicas of physical systems, facilitating the development of control algorithms and mission planning strategies.
4.1 4.1  Digital Twins (DTs)
@@
A digital twin refers to a simulated replica of a real robot that possesses similar attributes and can mimic the motion of its real-world pair within a virtual environment [96]. The concept of DT for field robots was first presented by NASA, where the digital entity was defined as an integrated multi-physics, multi-scale, probabilistic simulation of a vehicle or system that mirrors the life of its real-world peer [97]. This project envisioned four applications of the DT technology including pre-mission simulation, real-time modification, monitoring, and onboard forensics for comprehensive analyses.
Although two terms – digital shadow and twin – are used interchangeably in literature, there is a clear distinction between them. A shadow is a passive digital replica that mirrors the real-time entity without active predictive modeling or feedback loops [98]. It typically supports one-way (i.e., real2sim) data flow, and is limited to basic monitoring. A true DT would include high-fidelity representations of vehicle dynamics, sensor fusion models, environmental interactions, and a bi-directional data pipeline enabling predictive simulation [99]. For instance, the simplified setup shown in Fig. 6 does not offer predictive simulation; however, the digital robot mimics the motion of the real ROV, thereby qualifying as its shadow rather than a twin.
@@
The use of DTs has been reported in different branches of robotics including aerial, manufacturing, surgical, space, and marine robots. Generally, the raw sensor data acquired from the robot is processed with physics engines and reconstructed as a high-fidelity DT model [96]. Considering the complex nature of subsea environments, Van et al. [105] identify the communication between the two entities as a fundamental component of a successful subsea DT system. The authors highlight three use cases for DT: (i) ROV launch and recovery training, (ii) failure prediction and remaining useful life estimation, and (iii) trust-aware DT development to reduce operator’s decision-making time in critical situations. However, other literature [106, 107] argue that DT systems do not require a fully reliable or accelerated communication mode. Instead, during a communication disruption, the predictive nature of a DT system can be utilized to create a realistic representation as if communication had remained uninterrupted. More recently, Adetunji et al. [108] integrate SLAM and path planning capabilities in a versatile DT system. Other applications of marine DT include inspection [109], assistive control [58], trajectory planning [106], etc.
4.2 4.2  Underwater Simulators
@@
A simulator is essential to work with DTs as well as to mimic the external environment parameters. However, given the complex fluid dynamics and other physicochemical variations, it is extremely difficult to design an accurate physics-based model of an underwater environment. Engineers have created aquatic environments with physics and game engines (e.g. Unity, Blender, WebGL3D, ODE) as backends and have integrated basic properties of the medium such as buoyancy, waves, current profile, etc.; examples include UUV Simulator [23] and UWSim [104] that support Gazebo plugin [110] as well. Independently designed virtual ROVs or DTs of real ROVs are rendered in these platforms to test their kinematic properties and interactions with the surroundings. Researchers have studied different underwater vehicles such as gliders [106, 111], ROVs [107, 23], AUVs [109], and humanoid robots [81] in these simulators. Yang et al. [106] investigate a glider trajectory in a deep virtual ocean. Their simulations show the effects of the current gradient across the water column on the glider’s trajectory, especially during the vertical dives. Hu et al. [111] model a blended-wing-body underwater glider and investigate the accuracy of a pitch controller under ocean current influence.
Table 3 compares SOTA underwater simulators developed across academia and industry. Among proprietary simulators, Abyssal [101] and ACSM [112] offer a rich collection of underwater infrastructures and work-class ROV models, designed to train operators in mining, repairing, and maintenance of offshore rigs. PaleBlue [100] and VMAX Simulator [102] allow users to import custom-designed ROVs from CAD models and simulate their behavior, forces, frictions, etc. in the virtual world. Era Marine and EIVA have partnered with VideoRay to create advanced simulators for the VideoRay Pro5 and Defender ROVs. These simulators provide advanced tools for planning, surveying, and underwater inspections. M8-R Sim [113], originally developed for the MATE ROV Competition, has over $140140$ simulation scenarios, including those for monitoring the growth of Crown of Thorns Starfish (COTS). Unlike higher-end professional simulators, open-source alternatives offer fewer built-in scenarios and ROV models but allow for better customization and development according to mission-specific needs. For instance, UWSim [104], StoneFish [114], and HoloOcean [103] support ROS with sensor plugins e.g., side scan sonar, imaging sonar, acoustic beacon, DVL, magnetometer, force sensor, etc. UUV Simulator [23] integrates chemical particles and a simulation scenario to measure their concentration using a CPC (condensation particle counter) sensor. To summarize, open-source solutions are better suited for research and development with ROS support, whereas licensed products offer superior training opportunities for commercial ROV operation.
@@
Table 3: Comparison of underwater simulators are shown. Basic state estimation sensor plugins (e.g., IMU, GPS, camera, depth) are common for all simulators, therefore omitted here; : digital twin of a real ROV, other notations are explained in Sec. 2.
@@
  Sensor Plugin Built-in Scenarios Robot Models Optimal Usage
  PaleBlue [100] Laser Reefs, oil and gas rigs Triton XLX  VR-supported rig inspection
Abyssal [101] - Rigs, sand cloud, pipelines Abyssal ROV Manipulator control
M8-R Sim [113] - 140+ coral reefs with starfishes M8-R ROV Education & novice training
Era VideoRay [115] Sonar, lights 6 scenarios with shipwrecks VideoRay Pro5 , Defender  Training for VideoRay ROVs
GRi VROV [44] Sonar Pipelines and rigs LittleGeek, BigGeek Maintenance, rescue, mine detection
UUV Simulator [23] Sonar, DVL, CPC Shipwrecks, coast, BOP panels A9 , LAUV , Saga , RexROV Perception & control study
UUVSim [116] Sonar Open water, current, seabed 6T UUV, RexROV Hydrodynamic study, DL training
HoloOcean [103] Sonar, beacon, DVL Dams, pier harbors, pipelines 6 ROV agents Sim2Real, multi-agent control
UWSim [104] DVL, multi-beam sonar Dredging, pipelines, amphora Girona500  Planning, control, HIL study
UNav-Sim [117] LiDAR Algae, sandy pipelines BlueROV2 heavy  Navigation & inspection study
MARUS [118] LiDAR, sonar, DVL, AIS Ships, human divers, seabed One AUV Acoustic sensing study
StoneFish [114] Sonar, DVL Sea surface, seabed, wind, sun Girona500  Localization & navigation study
NaviSuite [36] Sonar, DVL, USBL Cables, pipelines, and ship anchors VideoRay Defender  Inspection, intervention
VMAX Simulator [102] Sonar Cables, ship with crane Triton , Comanche , Mojave  Manipulator control
ACSM [112] - Rigs and pipelines Triton XLX  Deployment & manipulator control
Unicom [119] - Rigs with valves, knobs, doors 2 work class ROVs Manuevering risk assessment
@@
5 5. Human-Machine Shared Autonomy
@@
Subsea robotic systems are categorized into three types according to their intelligent decision-making capabilities: (i) direct teleoperator control with no intelligence, (ii) shared/collaborative reasoning, and (iii) autonomous/self-sufficient systems. In direct teleoperation, the operator has total control over the robot’s motion and actions, generally through a live feed and a motion controller. However, these methods are constrained by communication latency and operator workload, making them less efficient for long-term complex tasks. Shared autonomy features bridge the gap between the human and the machine by allowing robots to perform certain tasks autonomously while the operator manages high-level decisions [120, 22]. This reduces the operator’s cognitive load, compensates for communication delays, and enhances operation safety. In contrast, AUVs function independently with minimal human intervention [121]. These robots rely on advanced algorithms for perception, decision-making, and execution, making them ideal for repetitive tasks in familiar environments. However, their utility is constrained by environmental uncertainties, the cost of logistics, and the limitations of current AI-based technologies [122].
5.1 5.1  Direct Teleoperation Mode
@@
Direct or hands-on teleoperation involves controlling the movements of an underwater robot using low-level commands without intelligent decision-making by the robot. The advantage of such systems is that the operator can execute instant decisions during the operation. It allows for responsive control, enabling precise adjustments that automated systems may struggle with. Despite high cognitive demand, direct teleoperation remains the preferred approach for manipulation tasks involving high risk and fine precision, where the operator controls one joint angle at a time in a joint-by-joint fashion [123]; examples include structure inspection [124], valve control [125], benthic sampling [123], etc.
@@
These systems are challenged by the inherent communication delays in underwater environments, which can reduce control precision and responsiveness. To address this, researchers have developed adaptive control strategies to compensate for both model uncertainties and external disturbances during teleoperation [125]. Furthermore, electromagnetic and optical communication modes offer promising solutions for establishing high-speed reliable links [126], improving the real-time commands and feedback flow in direct teleoperation systems.
@@
Table 4: The table summarizes various types of intelligent machines and their autonomous task-sharing capabilities reported across the subsea telerobotics literature. Criteria are defined in Sec. 5.2, notations are explained in Sec. 2.
@@
collaboration leader ROV follower ROVs search and rescue [6]
@@
demonstration relevant tasks adapt to novel tasks sample collection [134]
@@
Context-aware Provide contextual Medium reason a context, Inspection, surveillance, [5, 124]
exploration cues or guidance plan accordingly environmental monitoring
5.2 5.2   Human-Robot Collaborative Reasoning
@@
The goal of collaborative decision-making systems is to offer a blend of human oversight along with automated functions [135]. These systems allow operators to set high-level objectives while the robot handles low-level operations, thus reducing operator’s cognitive load and enabling more efficient operation [22]. For instance, an operator can issue hovering command; the robot determines its state using an IMU, DVL, and depth sensors to correct for any drift motion to remain stationary. More advanced reasonings are achieved in a supervisory manner (for a single robot) or with a leader-follower strategy (for a group of robots). Table 4 aims to present a structured overview of current capabilities and highlight how autonomy is distributed in SOTA subsea HMI systems. The clustering is motivated by a need to understand and compare the evolving roles of humans and machines in underwater missions. In modern telerobotic systems, autonomy is no longer binary (teleop vs. autonomous) but exists on a spectrum. Thus, the table categorizes task-sharing models by considering: (i) the nature of human intervention, (ii) the degree of machine intelligence, and (iii) the operational context (e.g., planning, fault handling, coordination). The cognitive load classification is informed by findings from user studies and qualitative insights reported by other researchers.
@@
For instance, mission planning, online mission adjustment, and fault diagnosis systems benefit from shared autonomy while keeping the operator in the loop [136], but without demanding continuous oversight. Research in supervised planning and navigation demonstrates that a new mission path suggested by a circumstance-aware robot is often safer and reduces task completion time [128, 31, 137]. The GALILEO planning interface adjusts the course of actions and resource allocations based on high-level goals set by the operator [138]. Additionally, intelligent robots are capable of detecting and reporting internal faults and data anomalies [129], reducing the need for constant operator oversight. Sharing scientific objectives with a context-aware robot is beneficial too, as the robot can search for specific cues or features under operator guidance. The CView project [5] provides an example; the operator defines the object of interest (e.g., wall, vessel, sluice) and the features to detect (e.g., cracks), enabling the robot to survey the objects and report key findings.
@@
SUPERMAN, the first supervisory subsea teleoperation interface, was proposed by NASA in $19781978$ [139]; however, the development has been slow due to the difficulty of designing high-level autonomy considering the environmental challenges. SAUVIM (1998199819981998[140] is one of the earliest semi-autonomous vehicles with a 6-DOF manipulator arm developed for deep water intervention tasks. More recently, supervisory control frameworks such as DexROV [79] and SHARC [24] have been developed where the operator performs remote manipulation tasks using language commands and hand gestures [128]. The GRASPER project [141] utilizes a pre-scanned scene point cloud to inform the robot about obstacle locations. The autonomous action is rendered in the UWSim simulator [104] in real-time, allowing the user to supervise grasping tasks in a virtual world.
@@
Moreover, contemporary researchers have developed adaptive supervision systems where the robot understands the operator’s overarching intention and adjusts the low-level reactions accordingly. Lee et al. [21] propose a shared teleoperation method where humans sketch a desired path on the visual console, guiding the intelligent system, which plans an optimum path along that drawn trajectory and executes the motion. The adaptive supervision system built within the GreenSea interface provides operators customized assistance based on their skill levels [120]. This helps novice pilots achieve performance comparable to that of experts.